OpenStudy (anonymous):

The temperature outside of John's house was negative seven degrees Fahrenheit. Throughout the day, it increased by nine degrees Fahrenheit. What was the final temperature outside?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@12rabbits

OpenStudy (anonymous):

\(-7+9=9-7\) is a start

OpenStudy (anonymous):

-7+9=2